OpenStudy (haltiamali):

Write a linear equation in slope-intercept form to model this situation. The population of Pitbulls is 6791 and it is decreasing at the rate of 7 per year.

OpenStudy (anonymous):

m=-7

OpenStudy (anonymous):

\[a _{n}=a+(n-1)d\]

OpenStudy (anonymous):

\[a _{n}=6791+(n-1)(-7)\]

OpenStudy (anonymous):

\[a _{n}=-7n+6798\]

OpenStudy (haltiamali):

thanks.

OpenStudy (anonymous):

hence \[f(x)=-7(x)+6798\]